A 1000 kg Civic is travelling at 30 m/s and accelerates to 40 m/s in 10 seconds.
a. What is the change in momentum?
b. What is the impulse?
c. What is the net force that causes the change?
1
Expert's answer
2020-12-14T04:43:26-0500

a. "\Delta p=mv_2-mv_1=m(v_2-v_1)=1000\cdot(40-30)=10000(kg\cdot m/s)"


b. "I=F\Delta t=1000\cdot\frac{40-30}{10}\cdot10=10000(kg\cdot m/s)"


c. "F=ma=1000\cdot\frac{40-30}{10}=1000(N)"